Given Ca=20 c=MPC=3/4 Ia=I=20 detemine equilibrium level of income when there is no government sector b; define equilibrium level of income when government spending is 25 &no taxation

1
Expert's answer
2022-02-04T09:20:10-0500

Qn. a

Equilibrium level of income without government sector is given as;

"Y=C+I"

"C=Ca+c\times Yd"

"c=\frac{3}{4}=0.75" "Ia=I=20"

but "Yd=Y-T"

"C=20+0.75Yd" but in this case there is no taxation so "Yd=Y"

"Y=(20+0.75Y)+20"

"Y=40+0.75Y"

"0.25Y=40"

"\bold{Y=160}" equilibrium income without government.


Qn. b

Equilibrium income with government spending of 25 is given as;

"Y=C+I+G"

but we have "Y=C+I\to 160" and the value of "G=25"

"Y=160+25\to 185"